OpenStudy (anonymous):

Choose the correct simplification of the expression (a^4/b^2)3

OpenStudy (zehanz):

I have nothing to choose here, so I'll just try simplifying it myself ;)\[\left( \frac{ a^4 }{ b^2 } \right)^3=\frac{ (a^4)^3 }{ (b^2)^3 }=\frac{ a ^{4*3} }{ b ^{2*3} }=\frac{ a ^{12} }{ b^6 }\]